How many molecules are there in 311g of NaCl


1
Expert's answer
2021-04-25T03:49:41-0400

Molar Mass of NaCl = 58.44

= 311/58.44 = 5.322 moles

1 mole = 6.022 × 1023

= 5.322 × 6.022 × 1023

= 3.205 × 1024 molecules